OpenStudy (anonymous):

How many terms are in an arith. sequence if the 1st term is 100 and the last term is -14 and d= -2

OpenStudy (jhannybean):

Arithmetic Progression\[\large a_{n} = a_{1}+(n-1)d\]\[\large a_{1} = 100 \]\[\large a_{n} = -14\]\[d=-2\]\[\large a_{-14} = 100 +(-14-1)(-2)\]\[\large a_{-14}= 100 -15(-2)\] I believe.